8-x = 6(x+1)
-- ------
3 7

Answer by stanbon(75887)

8-x = 6(x+1)
-- ------
3 7
--------------------------
Cross-multiply:
7(8-x) = 3*6(x+1)
56 - 7x = 18x + 18
25x = 58
x = 58/25
==================
Cheers,
Stan H.
==================